Suppose you’re given there is a memory access timeof 152 nanoseconds and an average page-faultservice time of 15 milliseconds.a.If one access out of 2000causes a page fault, what will be the Effective Access Time?b.How does this compare to main memory access time?
Answer #1

Given memory access time = 152 nano seconds

given average page fault-service time = 15 milliseconds.

Given one access out of 2000causes a page fault i.e,

page fault rate =1/2000=0.0005

Effective memory access time = (1-0.0005)*(152*10^-9) + (0.0005)*(15*10^-2)

                                                 =0.000076515

                                                = 7.65 micro seconds

B)if the effective memory time (7.65 microseconds) compared to memory access time(152 nano seconds) then effective memory time is 50.32 times greater than memory access time